# Stable-Pose: Leveraging Transformers for Pose-Guided Text-to-Image Generation

### v1

@@ Abstract
Controllable text-to-image (T2I) diffusion models have shown impressive performance in generating high-quality visual content through the incorporation of various conditions. Current methods, however, exhibit limited performance when guided by skeleton human poses, especially in complex pose conditions such as side or rear perspectives of human figures. To address this issue, we present Stable-Pose, a novel adapter model that introduces a coarse-to-fine attention masking strategy into a vision Transformer (ViT) to gain accurate pose guidance for T2I models. Stable-Pose is designed to adeptly handle pose conditions within pre-trained Stable Diffusion, providing a refined and efficient way of aligning pose representation during image synthesis. We leverage the query-key self-attention mechanism of ViTs to explore the interconnections among different anatomical parts in human pose skeletons. Masked pose images are used to smoothly refine the attention maps based on target pose-related features in a hierarchical manner, transitioning from coarse to fine levels. Additionally, our loss function is formulated to allocate increased emphasis to the pose region, thereby augmenting the model’s precision in capturing intricate pose details. We assessed the performance of Stable-Pose across five public datasets under a wide range of indoor and outdoor human pose scenarios. Stable-Pose achieved an AP score of 57.1 in the LAION-Human dataset, marking around 13% improvement over the established technique ControlNet. The project link and code is available at https://github.com/ai-med/StablePose.
@@ 1 Introduction
  • Introducing a hierarchical integration of pose masks for coarse-to-fine guidance, with a novel pose-masked self-attention mechanism and pose-mask guided loss function to enhance pose control.
@@ 3 Proposed Method
As shown in Figure 2, the denoising network $\epsilon_{\theta}$ adopts a UNet backbone, which is equipped with Stable-Pose for augmenting the latent encoding given the input conditional pose image. Let $\mathbf{\epsilon}_{\theta}(\mathbf{z}_{t},t,p)$, $t\in\{1,\cdots,T\}$, represent a T-step denoising UNet with gradients $\nabla{\theta}$ over a batch and input text prompt $p$. The conditional LDM is learned through $T$ steps by minimizing $\mathcal{L}=\mathop{\mathbb{E}}_{\mathbf{z},p,\varphi,\mathbf{\epsilon}\sim\mathcal{N}(0,I),t}\left[||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,\tau_{\theta}\left(p\right),\upsilon_{\theta}\left(\mathbf{z}_{t},\varphi\right)\right)||_{2}^{2}\right]$, where Stable-Pose $\upsilon_{\theta}$ conditions the latent encoding $\mathbf{z}_{t}$ on the input pose skeleton $\varphi\in\mathcal{R}^{h\times w\times 3}$, and $\tau_{\theta}$ is a text encoder that maps the text prompt $p$ to an intermediate sequence. The proposed framework is detailed in Figure 3. Stable-Pose aims at improving the frozen decoder in the UNet of SD to condition the input latent encoding $\mathbf{z}_{t}$ on the conditional pose image $\varphi$:
@@ 4 Experimental Results
Implementation Details. Similar to previous work Zhang et al. [2023], Mou et al. [2023], Zhao et al. [2024], we fine-tuned our model on SD with version 1.5. We utilized Adam Kingma and Ba [2014] optimizer with a learning rate of $1\times 10^{-5}$. For our proposed PMSA ViT module, we adopted a depth of 2 and a patch size of 2, where coarse-to-fine pose masks were generated using two Gaussian filters, each with a sigma value of 3 but with differing kernel sizes of 23 and 13, respectively. We will explore the effects of these hyperparameters with more details in Sec. 4.2. In the pose-mask guided loss function, we set an $\alpha$ of 5 as the guidance factor. We also followed Zhang et al. [2023] to randomly replace text prompts as empty strings at a probability of 0.5, which aims to strengthen the control of the pose input. During inference, no text prompts were removed and a DDIM sampler Song et al. [2020] with time steps 50 was utilized to generate images. On the Human-Art dataset, we trained all techniques, including ours for 10 epochs to ensure a fair comparison. On the LAION-Human subset, we trained Stable-Pose, HumanSD Ju et al. [2023b], GLIGEN Li et al. [2023] and Uni-ControlNet Zhao et al. [2024] for 10 epochs, while we used released checkpoints from other techniques due to computational limitations. The training was executed using two NVIDIA A100 GPUs, with our method completing in 15 hours for the Human-Art dataset and 70 hours for the LAION-Human subset. This represents a substantial decrease in GPU hours compared to SOTA techniques. For instance, the T2I-Adapter requires approximately 300 GPU hours to train on a large-scale dataset. In contrast, our approach requires less than a quarter of that time and still delivers superior performance.
@@ 4.1 Results
Results on the Outdoor and Indoor Poses. We extend the evaluation on both outdoor and indoor pose-guided T2I generation. We selected approximately 2,000 frames from the DAVIS dataset Perazzi et al. [2016], which comprises videos of human outdoor activities, as our outdoor pose assessment. In addition, we randomly chose around 2,000 images from the Dance Track dataset Sun et al. [2022], which is characterized by its group dance videos where most videos were shot indoors with multiple individuals and complex poses, as indoor pose-alignment evaluation. As shown in Table 3, the consistently highest AP and CAP scores achieved by Stable-Pose demonstrate its robustness in pose-controlled T2I generation across diverse environments, highlighting its potential as a backbone for pose-guided video generation.
@@ 4.2 Ablation Study
We conducted a comprehensive ablation study of Stable-Pose on the Human-Art dataset, including the effectiveness of pose masks, coarse-to-fine design, pose-mask guidance strength, and the effectiveness of PMSA and its ViT backbone.
@@
Coarse-to-Fine Masking Guidance. The granularity of pose-masks is specified by the Gaussian kernels in Gaussian Filters, where a larger kernel generates coarser pose-masks. We compared the results of constant granularity, fine-to-coarse as well as coarse-to-fine setting. All experiments are based on a ViT with PMSA and depth of 2 and Gaussian Filters with fixed sigma $\sigma=3$. As indicated in Table 5, the coarse-to-fine approach consistently offers the best performance across metrics for pose alignment and image quality. This improvement is likely due to its progressive refinement from coarser to finer granularity in pose regions. By methodically narrowing the focus to more precise controllable areas, this strategy smoothly enhances the accuracy of pose adjustments and the overall quality of the generated images.
@@
Pose-mask Guidance Strength in Loss. In our proposed loss function in Eq. (7), the parameter $\alpha$, referred to as the pose-mask guidance strength, controls the intensity of penalization applied to the pose regions. We evaluated the impact of varying $\alpha$ from 1 to 10 on pose alignment and image quality, with the results presented in Figure 5. These results demonstrate that higher values of $\alpha$ typically enhance model performance by increasingly directing the focus toward the pose areas, while also maintaining a high quality of generated images.
@@ 5 Discussion and Conclusion
We introduced Stable-Pose, a novel adapter that leverages vision transformers with a coarse-to-fine pose-masked self-attention strategy, specifically designed to efficiently manage precise pose controls during T2I generation. Stable-Pose outperforms current controllable T2I generation methods across five distinct public datasets, demonstrating high generation robustness across diverse environments and various pose scenarios. Notably, in complex scenarios involving rare poses such as side or back views and multiple figures, Stable-Pose exhibits exceptional performance in both pose and visual fidelity. This can be attributed to its advanced capability in capturing long-range patch-wise relationships between different anatomical parts of human pose images through our intricate conditioning design. As a result, Stable-Pose holds significant potential in applications demanding high pose accuracy. One limitation of Stable-Pose is its slightly longer inference time, primarily due to the integration of self-attention mechanisms within the ViT. In addition, despite excelling in pose control, Stable-Pose has yet to be evaluated with other conditions such as edge maps. Nevertheless, its design allows for straightforward adaptation to various external conditions, suggesting high potential for diverse applications.
Broader Impacts: Stable-Pose’s excellent pose control makes it a valuable tool in creating diverse artworks, animations, movies, and sports training programs. Additionally, it can be a reliable tool in healthcare and rehabilitation for correcting posture and preventing patients from musculoskeletal issues. However, there also exist the risks for misuse in generating forged images or videos, which could pose negative social concerns. Addressing these issues requires a combination of technical safeguards, regulatory frameworks, and ethical guidelines.
@@ A.1 Datasets and Preprocessing
A.2 Evaluation Metrics

### v2

@@ Abstract
Controllable text-to-image (T2I) diffusion models have shown impressive performance in generating high-quality visual content through the incorporation of various conditions. Current methods, however, exhibit limited performance when guided by skeleton human poses, especially in complex pose conditions such as side or rear perspectives of human figures. To address this issue, we present Stable-Pose, a novel adapter model that introduces a coarse-to-fine attention masking strategy into a vision Transformer (ViT) to gain accurate pose guidance for T2I models. Stable-Pose is designed to adeptly handle pose conditions within pre-trained Stable Diffusion, providing a refined and efficient way of aligning pose representation during image synthesis. We leverage the query-key self-attention mechanism of ViTs to explore the interconnections among different anatomical parts in human pose skeletons. Masked pose images are used to smoothly refine the attention maps based on target pose-related features in a hierarchical manner, transitioning from coarse to fine levels. Additionally, our loss function is formulated to allocate increased emphasis to the pose region, thereby augmenting the model’s precision in capturing intricate pose details. We assessed the performance of Stable-Pose across five public datasets under a wide range of indoor and outdoor human pose scenarios. Stable-Pose achieved an AP score of 57.1 in the LAION-Human dataset, marking around 13% improvement over the established technique ControlNet. The project link and code are available at https://github.com/ai-med/StablePose.
@@ 1 Introduction
  • Introducing a hierarchical integration of pose masks for coarse-to-fine guidance, with a novel pose-masked self-attention mechanism and pose-mask guided loss function. Stable-Pose is designed as a lightweight adapter that can be easily integrated into any pre-trained T2I diffusion models to effectively enhance pose control.
@@ 3 Proposed Method
As shown in Figure 2, the denoising network $\epsilon_{\theta}$ adopts a UNet backbone, which is equipped with Stable-Pose for augmenting the latent encoding given the input conditional pose image.
Let $\mathbf{\epsilon}_{\theta}(\mathbf{z}_{t},t,p)$, $t\in\{1,\cdots,T\}$, represent a T-step denoising UNet with gradients $\nabla{\theta}$ over a batch and input text prompt $p$. The conditional LDM is learned through $T$ steps by minimizing $\mathcal{L}=\mathop{\mathbb{E}}_{\mathbf{z},p,\varphi,\mathbf{\epsilon}\sim\mathcal{N}(0,I),t}\left[||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,\tau_{\theta}\left(p\right),\upsilon_{\theta}\left(\mathbf{z}_{t},\varphi\right)\right)||_{2}^{2}\right]$, where Stable-Pose $\upsilon_{\theta}$ conditions the latent encoding $\mathbf{z}_{t}$ on the input pose skeleton $\varphi\in\mathcal{R}^{h\times w\times 3}$, and $\tau_{\theta}$ is a text encoder that maps the text prompt $p$ to an intermediate sequence. The proposed framework is detailed in Figure 3. Stable-Pose aims at improving the frozen decoder in the UNet of SD to condition the input latent encoding $\mathbf{z}_{t}$ on the conditional pose image $\varphi$:
@@ 4 Experimental Results
Implementation Details. Similar to previous work Zhang et al. [2023], Mou et al. [2023], Zhao et al. [2024], we fine-tuned our model on SD with version 1.5. We utilized Adam Kingma and Ba [2014] optimizer with a learning rate of $1\times 10^{-5}$. For our proposed PMSA ViT module, we adopted a depth of 2 and a patch size of 2, where coarse-to-fine pose masks were generated using two Gaussian filters, each with a sigma value of 3 but with differing kernel sizes of 23 and 13, respectively. We will explore the effects of these hyperparameters with more details in Sec. 4.2. In the pose-mask guided loss function, we set an $\alpha$ of 5 as the guidance factor. We also followed Zhang et al. [2023] to randomly replace text prompts as empty strings at a probability of 0.5, which aims to strengthen the control of the pose input. During inference, no text prompts were removed and a DDIM sampler Song et al. [2020] with time steps 50 was utilized to generate images. On the Human-Art dataset, we trained all techniques, including ours for 10 epochs to ensure a fair comparison. On the LAION-Human subset, we trained Stable-Pose, HumanSD Ju et al. [2023b], GLIGEN Li et al. [2023] and Uni-ControlNet Zhao et al. [2024] for 10 epochs, while we used released checkpoints from other techniques due to computational limitations. The training was executed using two NVIDIA A100 GPUs, with our method completing in 15 hours for the Human-Art dataset and 70 hours for the LAION-Human subset. This represents a substantial decrease in GPU hours compared to SOTA techniques. For instance, the T2I-Adapter requires approximately 300 GPU hours to train on a large-scale dataset. In contrast, our approach requires less than a quarter of that time and still delivers superior performance. We kept the same seed list for all techniques, including ours, during both training and inference time, to ensure fair comparison and reproducibility. More detailed information is provided in Sec. A.2.
@@ 4.1 Results
Stable-Pose as a generic adapter. Stable-Pose is designed as a lightweight generic adapter that can be easily integrated into any pre-trained T2I diffusion models to effectively enhance pose control. To further validate its generalizability, we conducted additional experiments by applying Stable-Pose on top of a pre-trained HumanSD Ju et al. [2023b] model. As shown in Table 2, the inclusion of Stable-Pose considerably improved the baseline HumanSD by over 10% in AP and 12% in KID, highlighting its high generalizability and effectiveness in enhancing both pose control and image quality.
Table 2: Stable-Pose as a lightweight adapter on pre-trained HumanSD model.
HumanSD 44.57 69.68 1.37 10.03 2.70 32.24
Stable-Pose 48.88 70.83 1.50 11.12 2.35 32.60
HumanSD+Stable-Pose 49.24 71.01 1.42 10.42 2.37 32.16
@@
Results on the Outdoor and Indoor Poses. We extend the evaluation on both outdoor and indoor pose-guided T2I generation. We selected approximately 2,000 frames from the DAVIS dataset Perazzi et al. [2016], which comprises videos of human outdoor activities, as our outdoor pose assessment. In addition, we randomly chose around 2,000 images from the Dance Track dataset Sun et al. [2022], which is characterized by its group dance videos where most videos were shot indoors with multiple individuals and complex poses, as indoor pose-alignment evaluation. As shown in Table 4, the consistently highest AP and CAP scores achieved by Stable-Pose demonstrate its robustness in pose-controlled T2I generation across diverse environments, highlighting its potential as a backbone for pose-guided video generation. Further results can be found in Table A.6.
@@ 4.2 Ablation Study
We conducted a comprehensive ablation study of Stable-Pose on the Human-Art dataset, including the effectiveness of pose masks, coarse-to-fine design, pose-mask guidance strength, and the effectiveness of PMSA and its ViT backbone. Further ablations on model parameters can be found in Sec. A.6.
@@
Coarse-to-Fine Masking Guidance. The granularity of pose-masks is specified by the Gaussian kernels in Gaussian Filters, where a larger kernel generates coarser pose-masks. We compared the results of constant granularity, fine-to-coarse as well as coarse-to-fine setting. All experiments are based on a ViT with PMSA and depth of 2 and Gaussian Filters with fixed sigma $\sigma=3$. Details can be found in Sec. A.3. As indicated in Table 6, the coarse-to-fine approach consistently offers the best performance across metrics for pose alignment and image quality. This improvement is likely due to its progressive refinement from coarser to finer granularity in pose regions. By methodically narrowing the focus to more precise controllable areas, this strategy smoothly enhances the accuracy of pose adjustments and the overall quality of the generated images.
@@
Pose Encoding in Stable-Pose. To further validate the design of pose encoding in Stable-Pose, we implemented an ablation study by removing either the pose encoder $\beta$ or PMSA-ViT, retaining only one type of pose encoding. The results in Table 8 show that using only PMSA-ViT yields an AP of 36.48, which is expected due to the absence of color-coding information for distinguishing body parts. While using $\beta$ alone increases the AP to 45.03. However, the most significant improvement is observed when integrating both local and global information encoding into the Stable-Pose architecture, achieving the highest AP of 48.88.
Table 8: Ablation study on the pose encoding design in Stable-Pose.
w/o $\beta$ Encoder 36.48 68.91 1.55 11.17 2.76 31.90
w/o PMSA-ViT Encoder 45.03 70.38 1.52 13.67 2.49 32.53
w/ both $\beta$ PMSA-ViT Encoder 48.88 70.83 1.50 11.12 2.35 32.60
Pose Masks During Inference. We incorporate the pose masks during inference by default to enhance pose control. To further validate their effectiveness, we additionally conducted experiments with removing the pose masks during inference. As shown in Table LABEL:tab:ablation_posemasks_inference, this led to approximately a 3% drop in AP. This could be due to two main reasons: 1) The pose masks provided additional guidance, thus enhancing control; 2) The inclusion of pose masks maintains consistency between the model’s behavior during training and inference. Thus, including pose masks benefits pose control in the generation.
Table 9: Ablation study on the presence of pose masks during inference.
w/o mask in inference 45.93 70.51 1.52 13.11 2.55 32.68
w/ mask in inference 48.88 70.83 1.50 11.12 2.35 32.60
@@
Pose-mask Guidance Strength in Loss. In our proposed loss function in Eq. (7), the parameter $\alpha$, referred to as the pose-mask guidance strength, controls the intensity of penalization applied to the pose regions. We evaluated the impact of varying $\alpha$ from 1 to 10 on pose alignment and image quality, with the results presented in Figure 5. Increasing $\alpha$ in our proposed loss means putting more attention on the foreground pose regions. However, when $\alpha$ is getting too large, it forces the model to learn irrelevant texture information like clothing, which negatively impacts training and slightly decreases AP. Despite this, Stable-Pose still outperforms others across an $\alpha$ range of 1-10. Notably, increasing $\alpha$ has a significant impact on FID, worsening it from 11.0 at $\alpha$=5 to 13.0 at $\alpha$=10. This indicates that focusing solely on the pose regions may decrease the quality of generated content in non-pose regions. Thus there exists a slight trade-off in selecting $\alpha$ to maintain both high pose accuracy and image quality, in which a value around 5 to 6 turns out to be optimal.
@@ 5 Discussion and Conclusion
We introduced Stable-Pose, a novel adapter that leverages vision transformers with a coarse-to-fine pose-masked self-attention strategy, specifically designed to efficiently manage precise pose controls during T2I generation. Stable-Pose outperforms current controllable T2I generation methods across five distinct public datasets, demonstrating high generation robustness across diverse environments and various pose scenarios. Notably, in complex scenarios involving rare poses such as side or back views and multiple figures, Stable-Pose exhibits exceptional performance in both pose and visual fidelity. This can be attributed to its advanced capability in capturing long-range patch-wise relationships between different anatomical parts of human pose images through our intricate conditioning design. As a result, Stable-Pose holds significant potential in applications demanding high pose accuracy. It can be easily integrated into any pre-trained T2I diffusion models as a lightweight generic adapter to effectively enhance pose control. One limitation of Stable-Pose is its slightly longer inference time (Sec. A.2), primarily due to the integration of self-attention mechanisms within the ViT. In addition, despite excelling in pose control, Stable-Pose has yet to be evaluated with other conditions such as edge maps. Nevertheless, its design allows for straightforward adaptation to various external conditions, suggesting high potential for diverse applications.
Broader Impacts: Stable-Pose’s excellent pose control makes it a valuable tool in creating diverse artworks, animations, movies, and sports training programs. Additionally, it can be a reliable tool in healthcare and rehabilitation for correcting posture and preventing patients from musculoskeletal issues.
@@ A.1 Datasets and Preprocessing
A.2 Training and Evaluation
To ensure a fair comparison, we trained and evaluated all the techniques on Human-Art dataset. On LAION-Human subset, due to computational constraints, we only trained our method, HumanSD, GLIGEN and Uni-ControlNet, while ControlNet and T2I-Adapter were evaluated using the released checkpoints. The rationale behind this was that HumanSD had been previously trained on a segment of LAION-Human, which might overlap with our validation set, leading to potential data leakage. GLIGEN had been trained with considerably more GPU hours compared to other techniques, posing a risk of an unfair comparison. Uni-ControlNet’s available checkpoints are tailored for a multi-condition model, however, our research is dedicated exclusively to pose control, necessitating further training. Conversely, for ControlNet and T2I-Adapter, the amount of data and GPU hours for training were on par with our methods, allowing for fair comparison using the checkpoints they provided.
Training was conducted on two NVIDIA A100 GPUs. Our method required 15 hours to complete training on the Human-Art dataset and 70 hours for the LAION-Human subset. During training, the peak RAM usage for our technique was 25 GB. We reported the number of trainable parameters and training time on Human-Art dataset for each technique in Table LABEL:appendix::tab::params_inference_speed. Our technique has comparable number of trainable parameters and training time to ControlNet but produces significantly better results, highlighting the effectiveness of our model design. We also reported the inference speed of each technique in Table LABEL:appendix::tab::params_inference_speed, measured by the average seconds needed to generate an image. To ensure a fair comparison, these measurements were conducted on the same NVIDIA A100 GPUs. Due to the self-attention mechanism in the ViT, our method’s inference speed is slightly slower than that of ControlNet.
Table A.1: Trainable parameters (millions), training time on Human-Art (hours), and inference time (seconds per image) for each technique.
Method T2I-Adapter ControlNet Uni-ControlNet GLIGEN HumanSD Stable-Pose
Train. params. 77 361 411 209 859 364
Train. time 12.9 14.0 14.2 20.2 19.6 14.3
Infer. time 4.45 5.76 6.20 9.48 4.20 6.37
A.3 Selecting Gaussian Filters in PMSA
A common way to choose the kernel size $k$ of a Gaussian Filter with sigma $\sigma$ is $k=2\times\lceil 3\sigma\rceil+1$. This ensures that the kernel captures the majority of the Gaussian function’s weight. When using the kernels and sigmas as specified in the above equation, we observed that the difference between the generated coarse and fine masks was too large to yield satisfactory results. We measured this via the computation of the activation rate, reported in Table A.2. We applied the pose mask to the patches and then calculated the ratio of pose-masked patches to the total number of patches. We opted for an alternative method to control the blurring effect by fixing $\sigma$ and adjusting $k$, which changes the number of pixels to which the Gaussian weights are applied. Larger values for $k$ dilate the mask, resulting in coarser masks.
In our experiment, we fix $\sigma$ at 3 and then change $k$ for obtaining different kernels. The activation rate of patches under the fixed $\sigma$ and varying $k$ is reported in Table A.3, where $k=23$ is the largest kernel size for $\sigma=3$, indicating that any larger $k$ does not affect the activation rate. This size was chosen to generate the most coarse mask. To ensure smooth guidance of our pose-mask in the proposed PMSA, we selected kernel sizes with approximately one percent difference in activation rate. As a result, $k=13$ was chosen to generate the fine mask. For our ablation study with 3 pose masks, we selected $k=9$, which also yielded about a one percent difference in activation rate compared to $k=13$. Details of this ablation study can be found in Sec. A.6.
Table A.2: Ratios of unmasked patches (%) under different sigma $\sigma$ and kernel size $k$ in a coarse-to-fine manner.
$\sigma=4,k=25$ $\sigma=3,k=19$ $\sigma=2,k=13$ $\sigma=1,k=7$
15.83 14.12 12.38 10.48
Table A.3: Ratios of unmasked patches (%) under fixed sigma $\sigma=3$ and varying kernel size $k$ in a coarse-to-fine manner. Selected ones are in bold.
23 21 19 17 15 13 11 9 7 5 3 1
14.26 14.17 14.12 13.98 13.77 13.21 12.78 12.14 11.53 10.81 10.06 9.09
A.4 Evaluation Metrics
@@
Challenges in the Current Metrics. We followed the common metrics widely adopted in the current generative AI field, however, despite the rapid advancements in generative AI, existing metrics have not evolved to provide a more accurate evaluation Jayasumana et al. [2024]. Some of the issues are 1) CLIP score relies on cosine similarity between the model’s semantic understanding and the given text, which may not align with pose assessments or the relevance of generated images. Additionally, this score is sensitive to the arrangement and composition of elements in images; even minor changes can result in significant fluctuations in the score, which may not accurately reflect the overall generative quality. Ruiz et al. [2023] suggests a Dino-based score; 2) FID estimates the distance between a distribution of Inception-v3 features of real images and those of images generated by the generative models. However, Inception’s poor representation of the rich and varied content generated by modern text-to-image models incorrect normality assumptions and poor sample complexity Jayasumana et al. [2024]. Thus, the FID score does not account for semantic correctness or content relevance—specifically pose—in relation to the specified text or conditions. Relying solely on FID and CLIP scores does not provide a comprehensive assessment of the generative model. Therefore, we further evaluated our method with a new state-of-the-art metric CMMD Jayasumana et al. [2024], which is based on richer CLIP embeddings and the maximum mean discrepancy distance with the Gaussian RBF kernel. It is an unbiased estimator that does not make any assumptions on the probability distribution of the embeddings, offering a more robust and reliable assessment of image quality. As shown in Table A.4, our method achieves better CMMD value compared to HumanSD, demonstrating comparably high image quality.
Table A.4: CMMD evaluation for HumanSD and Stable-Pose.
HumanSD Stable-Pose
CMMD $\downarrow$ 5.027 5.025
A.5 Detailed Results
We reported detailed results for the UBC Fashion, DAVIS, and Dance Track datasets in Tables A.5&A.6, where KID is multiplied by 100 for readability. The results demonstrate that our Stable-Pose consistently provides better controllability over poses, even for challenging poses such as back poses.
Table A.5: Detailed results on UBC Fashion dataset with front, side, and back orientations.
Orientation Method Pose Accuracy Image Quality T2I Alignment
Front T2I-Adapter 72.20 50.16 0.61 6.57 3.95 32.11
ControlNet 74.64 48.11 0.65 4.50 3.91 32.06
Uni-ControlNet 79.47 51.21 0.57 6.14 4.40 32.14
GLIGEN 73.97 50.95 0.44 8.94 4.49 30.35
HumanSD 76.83 51.45 0.60 6.55 4.89 31.14
Stable-Pose (Ours) 87.26 51.90 0.56 6.05 4.57 32.05
Side T2I-Adapter 36.80 44.38 0.75 6.10 3.81 32.16
ControlNet 52.83 46.16 0.80 5.05 3.95 32.13
Uni-ControlNet 58.26 46.70 0.70 5.73 4.28 32.15
GLIGEN 45.32 45.68 0.69 6.58 3.94 30.75
HumanSD 57.09 45.85 0.78 5.14 4.75 31.95
Stable-Pose (Ours) 69.76 47.00 0.71 5.10 4.43 32.25
Back T2I-Adapter 6.03 22.38 0.91 8.49 4.22 32.20
ControlNet 23.68 24.61 0.97 7.53 4.01 32.15
Uni-ControlNet 19.97 23.79 0.93 7.20 4.37 32.19
GLIGEN 4.45 18.36 0.90 10.83 4.32 30.29
HumanSD 11.05 21.15 0.91 9.12 4.99 32.01
Stable-Pose (Ours) 29.08 25.21 0.93 6.24 4.56 32.11
Table A.6: Detailed results on DAVIS dataset and Dance Track dataset.
Dataset Method Pose Accuracy Image Quality T2I Alignment
DAVIS T2I-Adapter 20.28 60.76 1.90 3.69 2.19 31.05
ControlNet 30.13 60.81 1.67 3.81 2.21 30.72
Uni-ControlNet 37.64 60.57 1.56 3.60 1.83 30.63
GLIGEN 12.17 59.87 1.83 3.72 1.69 30.16
HumanSD 38.32 60.59 1.37 2.94 1.82 28.49
Stable-Pose (Ours) 42.87 62.43 1.45 3.54 1.84 30.56
Dance Track T2I-Adapter 10.36 72.38 4.82 19.76 5.64 33.01
ControlNet 16.45 73.16 5.78 20.92 5.13 32.76
Uni-ControlNet 25.22 73.62 4.79 21.89 4.75 32.83
GLIGEN 5.61 72.57 4.72 20.13 4.53 32.37
HumanSD 24.13 69.93 3.78 23.15 8.58 31.64
Stable-Pose (Ours) 28.58 74.97 4.81 17.28 4.74 32.90
A.6 Additional Ablation Studies
ViT Parameters. We studied the impact of different ViT parameters, including depth and patch sizes, as reported in Table A.7 and A.8. Table A.7 shows results for varying depths with a fixed patch size of 2, using coarse-to-fine pose masks as guidance. While these masks provide smooth guidance, a higher depth with overly fine masks might lose valuable information from previous attention layers. Table A.8 presents results for a ViT with PMSA of depth 2 and different patch sizes. Our default patch size of 2 yields the best results, likely because the latent encoding has typically lower dimension (here $64\times 64$) compared to the dimensions of input high-resolution images. Larger patches may dilute learning of interconnections among anatomical parts for encompassing too much information.
Table A.7: Comparison of different depths in our proposed ViT. Kernel sizes of Gaussian Filters are noted.
Depth AP $\uparrow$ CAP $\uparrow$ PCE $\downarrow$ FID $\downarrow$ KID $\downarrow$ CLIP-score $\uparrow$
1 (23) 47.96 70.75 1.57 11.94 2.45 32.55
2 (23, 13) 48.88 70.83 1.50 11.12 2.35 32.60
3 (23, 13, 9) 47.64 71.13 1.55 12.22 2.65 32.58
Table A.8: Comparison of different patch sizes in our proposed ViT.
Patch size AP $\uparrow$ CAP $\uparrow$ PCE $\downarrow$ FID $\downarrow$ KID $\downarrow$ CLIP-score $\uparrow$
2 48.88 70.83 1.50 11.12 2.35 32.60
4 47.51 70.85 1.53 13.39 2.49 32.58
8 45.69 70.44 1.59 13.14 2.47 32.64
A.7 Attention Maps of PMSA
In Figure A.2, we present examples of attention maps extracted from our ViT. These maps are derived from the last block of the ViT and represent an average across all attention heads. The attention maps were approximately overlapped with the pose region, which corroborates the objectives of our proposed PMSA. This alignment underscores the efficacy of our model in focusing on relevant pose features, a critical aspect of our approach to improving model interpretability.
Figure A.2: Samples from LAION-Human dataset with original images and attention maps.
A.8 Failure Case
Stable-Pose may face failure cases when generating the wrong number of people in very crowded scenes. Stable-Pose enhances SD’s accuracy in pose-mask regions, whereas a pre-trained SD may still produce human-like figures in the background. For example, in the first row of Figure A.3, Stable-Pose generates an extra half-shaped person on the very right side.
Figure A.3: Failure cases sampled from Dance Track dataset. Unexpected humans generated are marked in red circles.
A.9 Extreme cases
The generation of some extreme pose cases is shown in Figure A.4, such as bending the upper body backward in some dancing poses. As shown in the figures, Stable-Pose still maintains very high pose accuracy on generated images under these challenging scenarios, whereas ControlNet fails to depict the correct pose and body structures.
Figure A.4: Extreme poses sampled from the Human-Art dataset. ControlNet creates wrong limbs in extreme cases while Stable-Pose achieves accurate target poses.
A.10 More Visualization Results
We include additional qualitative results from the five datasets we evaluated in Figure A.5,A.6,A.7,A.8, & A.9.
Figure A.5: Results on Human-Art dataset.
Figure A.6: Results on LAION-Human dataset.
Figure A.7: Results on UBC Fashion dataset with side orientation.
Figure A.8: Results on UBC Fashion dataset with back orientation.
Figure A.9: Results on Dance Track (first row) and DAVIS dataset (the rest rows).